OpenStudy (anonymous):

describe how you know what number to add to both sides of an equation containing subtraction.

OpenStudy (anonymous):

Did they give you an example of an equation?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Well, In an example of : 2x = - 2x + 4 Which number would you add to both sides to cancel out?

OpenStudy (anonymous):

it depends on what number you want to move to the other side using @wmckinely 's example 2x = - 2x + 4 if we wanted to move the -2x on the right side to the left side we would have to negate it on the right side this could be done by adding 2x to both sides because -2x +2x =0 the result would look like 2x+2x = 4 now say we wanted to move the 2x on the left side to the right side again we need to negate that term on the left side and this can be done by adding -2x or subtracting 2x to both sides the result will look like 0=-4x +4 or if we wanted to move the 4 from the right side to the left side negating the 4 on the right side means we would have to subtract 4 to both sides because 4 -4 =0 and the result would look like -4 +2x=-2x

OpenStudy (anonymous):

i don't need the answer anymore but thanks :)